2. DRUG Q has an ED50 of 20m/kg in humans and the LD50 is 100mg/kg. DRUG Z has an ED50 of 10mg/kg, and the LD50 is 20mg/kg. Calculate the therapeutic index for both drugs (show all work!). Which drug would be easier for a person to overdose
Answer #1

DRUG-Q

ED50 (Effective dose) = 20mg/kg

LD50 (lethal dose) = 100mg/kg

So, Therapeutic index = LD50/ED50

= 100/20 = 5mg/kg

DRUG-Z

ED50 = 10mg/kg

LD50 = 20mg/kg

So, Therapeutic index = 20/10 = 2mg/kg

The therapeutic index of drug -Q is 5mg/kg. It means that one has to ingest 5 times the effective dose of a drug to ingest the lethal dose. Similarly, for drug -Z, one has to ingest only 2 times the effective dose of a drug to ingest its lethal dose. Therefore, therapeutic index is inversely proportional to the drug overdose. It means that drug-Q can be overdosed.

Hope it helps, Good luck !!!!!
Please do press like :)